# FedPeWS: Personalized Warmup via Subnetworks for Enhanced Heterogeneous Federated Learning

### v1

@@ Appendix C Experimental Results
Figure 11 displays the validation accuracy over $T=300$ communication rounds for both scenarios. The leftmost plot shows the accuracy of the global model, while the middle and rightmost plots the accuracy for each of the participants. Both mask length scenarios converge to a comparable accuracy levels, with a marginal difference of $0.5\%$ higher accuracy in the scenario where $|m_{1}|<|m_{2}|$. This is likely due to the larger mask size, which aids in learning the more complex CIFAR-10 dataset held by Participant 2.
C.3 Sensitivity Analysis
@@
C.4 Larger number of participants
Although our primary focus is on the cross-silo setting, we extend our study to include a large-scale scenario involving 200 participants on the CIFAR-MNIST dataset. We adopt a Dirichlet partition strategy with concentration parameter $\alpha=0.5$ and implement this scenario with a partial participation rate of $0.1$. The outcomes of this experiment, as depicted in Figure 12, indicate a superior performance compared to the conventional FedAvg algorithm, thereby further substantiating the validity and effectiveness of our proposed method. The parameters set for FedPeWS are: $\tau=25$ and $\lambda=0.5$.
@@ C.4 Larger number of participants
Appendix D Neuron Activations
@@
    After switching to full mask $(t>W)$: (i) there is a noticeable increase in activations for both participants upon switching to using full masks, (ii) Participant 1 with its originally initialized subnetwork 1, shows a substantial increase in activations compared to subnetwork 2 across all layers. The same pattern is obesrved for Participant 2 with subnetwork 2.
These findings suggest that the personalized warmup strategy helps the network learn which paths to follow when specific data points are fed into the network. This supports the superiority of our method and corroborates the claims made in the main paper.

### v2

@@ C.2 Comparison to FedProx
Figure 11 displays the validation accuracy over $T=300$ communication rounds for both scenarios. The leftmost plot shows the accuracy of the global model, while the middle and rightmost plots show the accuracy for each of the participants. Both mask length scenarios converge to comparable accuracy levels, with a marginal difference of $0.5\%$ higher accuracy in the scenario where $|m_{1}|<|m_{2}|$. This is likely due to the larger mask size, which aids in learning the more complex CIFAR-10 dataset held by Participant 2.
@@
C.5 Sensitivity Analysis
@@
C.6 Larger number of participants
C.6.1 $\mathbf{N=\{5,10,20\}}$ participants.
Figure 12 demonstrates the performance of our proposed algorithm on the CIFAR-10 and CIFAR-100 datasets [50] using a non-overlapping class splitting strategy, where an equal number of classes are sampled without replacement, totaling to the number of participants multiplied by the number of classes assigned to each participant. This represents a heterogeneous setting. For this comparison, we use FedProx with a scalar multiplier of $0.01$ as the baseline algorithm. The FedPeWS algorithm combines FedProx and our proposed method, referred to as FedProx+PeWS. As evident from the plots, our algorithm consistently outperforms the baseline, delivering a significant performance boost in heterogeneous scenarios and showcasing its effectiveness across various settings.
Figure 12: Visualization of global model performance with $N={5,10,20}$ participants on the CIFAR-10 and CIFAR-100 datasets. Note: NOC refers to non-overlapping classes partition; e.g. for the CIFAR-100 experiment, we randomly select 5 classes out of 100 for each of the 20 participants without replacement.
C.6.2 $\mathbf{N=200}$ participants.
Although our primary focus is on the cross-silo setting, we extend our study to include a large-scale scenario involving 200 participants on the CIFAR-MNIST dataset. We adopt a Dirichlet partition strategy with concentration parameter $\alpha=0.5$ and implement this scenario with a partial participation rate of $0.1$. The outcomes of this experiment, as depicted in Figure 13, indicate a superior performance compared to the conventional FedAvg algorithm, thereby further substantiating the validity and effectiveness of our proposed method. The parameters set for FedPeWS are: $W=25$ and $\lambda=0.5$.
@@ C.6 Larger number of participants
Appendix D Neuron Activations Study
@@
    After switching to full mask $(t>W)$: (i) there is a noticeable increase in activations for both participants upon switching to using full masks, (ii) Participant 1 with its originally initialized subnetwork 1, shows a substantial increase in activations compared to subnetwork 2 across all layers. The same pattern is observed for Participant 2 with subnetwork 2.
These findings suggest that the personalized warmup strategy helps the network learn which paths to follow when specific data points are fed into the network. This supports the superiority of our method and corroborates the claims made in the main paper. On the other hand, for the FedAvg algorithm (see Figure 15), we observe the absence of a similar activation pattern. Data points from different participants activate overlapping regions of the model, which results in slower convergence.
Figure 15: Neuron activation study on the Synthetic-32K dataset with a global learning rate $\eta_{g}=1.0$ using FedAvg algorithm.